OpenStudy (anonymous):

Let Yn be the largest order statistic in a sample of size n from the uniform distribution on [0, t]. Show that Yn converges in probability to t, that is, that P (|Yn - t| >=e) approaches to 0 as n approaches to infinity.

OpenStudy (anonymous):

There is so much context to this question that you should know and the typical math major hasn't committed to memory.

OpenStudy (anonymous):

but all of these answers for the exam :(

OpenStudy (anonymous):

What is a largest order statistic? What about this can you tell us?

OpenStudy (anonymous):

\[ \lim_{n\to \infty}\Pr(|Y_n-t|\geq \epsilon) = 0 \]

OpenStudy (anonymous):

p, (1-p)*p/n

OpenStudy (anonymous):

Do I have the formula right?

OpenStudy (anonymous):

So is \(Y_n\) the largest value in the sample?

OpenStudy (anonymous):

Start making sense of it. Don't have others do all the work for you.

OpenStudy (anonymous):

\[\lim_{n \rightarrow 0} P(|Y _{n} - \theta|\ge \epsilon) \rightarrow0\]

OpenStudy (anonymous):

\(n\) is approaching \(\infty\).

OpenStudy (anonymous):

yes Yn is the largest samle in a sample

OpenStudy (anonymous):

oh, yeah sorry n approaches to infinity

OpenStudy (anonymous):

So is \(\theta\) the largest possible sample point? While \(Y_n\) is the largest in the sample.

OpenStudy (anonymous):

i did some problems about distribution, z score, and others by myself, and got a right answers, but some of them are really very hard

OpenStudy (anonymous):

i couldnt understand the "converges in probability to"

OpenStudy (anonymous):

\[ \Pr(X\geq \epsilon) = 1-\Pr(X<\epsilon) \]

OpenStudy (anonymous):

This property may or may not help us.

OpenStudy (anonymous):

i think that may

OpenStudy (anonymous):

It really doesn't matter what "converges in probability to" means since they gave us the definition, but I think the idea is the theoretical value you'd expect after infinite trials.

OpenStudy (anonymous):

in any case thank u very much

OpenStudy (anonymous):

You have an interval: \[ [Y_n,t] \]After \(n\) trials.

OpenStudy (anonymous):

It's a uniform distribution, right?

OpenStudy (anonymous):

Actually I think I know how we can do this.

OpenStudy (anonymous):

\[(Y _{n}; \theta) \]

OpenStudy (anonymous):

If we do \(n\) trials, and have \(Y_n\), then the probability that the next trial is above \(Y_n\) is\[ \frac{\theta -Y_n}{\theta} \]

OpenStudy (anonymous):

yes right

OpenStudy (anonymous):

The probability we surpass it by the \(m\)th trial is \[ \left(\frac{Y_n}{\theta}\right)^m\left(\frac{\theta -Y_n}{\theta}\right) \]

OpenStudy (anonymous):

The probability we surpass it by the \(m\)th trial is \[ 1-\left(\frac{Y_n}{\theta}\right)^m \]

OpenStudy (anonymous):

why the m-th trial?

OpenStudy (anonymous):

We might not surpass it right after? I dunno.

OpenStudy (anonymous):

All I can really say is that The gap between \(Y_n\) and \(\theta\) is narrowing, but I'm not sure how we'll exactly prove it.

OpenStudy (anonymous):

The gap can't get any wider.

OpenStudy (anonymous):

Wait a minute! I think I really got it this time!

OpenStudy (anonymous):

you gave me some keys, i think i need more theory

OpenStudy (anonymous):

Should be \(n\to \infty\)

OpenStudy (anonymous):

but why n approaches to 0, not to infinity?

OpenStudy (anonymous):

Because I copied and pasted your mistake.

OpenStudy (anonymous):

oh, sorry :)

OpenStudy (anonymous):

yes, it' right, but they also can be equal

OpenStudy (anonymous):

Well it works for the \(|0|=0\) case as well.

OpenStudy (anonymous):

ok, thank you very much for your time

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."